OpenStudy (anonymous):

(x-5) divided by 3 plus x divided by 6 = (2x-1) divided by 8

OpenStudy (kendricklamar2014):

Step 1: Simplify both sides of the equation. 1/2x + -5/3 = 1/4x + -1/8

OpenStudy (kendricklamar2014):

Step 2: Subtract 1/4x from both sides. 1/2x + -5/3 - 1/4x = 1/4x + -1/8 - 1/4x

OpenStudy (kendricklamar2014):

which is: 1/4x + -5/3 = -1/8

OpenStudy (kendricklamar2014):

Step 3: Add 5/3 to both sides. 1/4x + -5/3 + 5/3 = -1/8 + 5/3